OpenStudy (lucaz):

what can I do to find the lim x->0 ((sqrt(x+4)-2)/x ?

myininaya (myininaya):

try rationalizing the numerator

OpenStudy (lucaz):

how?

myininaya (myininaya):

multiplying by top's conjugate do it to both top and bottom

OpenStudy (anonymous):

\[\frac{\sqrt{(x+4)} -2}{x}\] I would use L'H'S derive top and bottom

OpenStudy (anonymous):

( f(g)-f(h) ) ' = f(g)' - f(h) '

OpenStudy (anonymous):

never mind, I am not a good helper.

OpenStudy (lucaz):

@myininaya I got this

OpenStudy (lucaz):

1/4.. thanks

myininaya (myininaya):